OpenStudy (m):

Mx(t) = p^et/[1-q^et] dMx(t)/dt =?

OpenStudy (yuki):

what does p and q stand for ? could you use the equation button to express it ?

OpenStudy (dumbcow):

quotient rule \[\frac{dM}{dt}= \frac{(e ^{t}\ln p (p ^{e ^{t}})(1-q ^{e ^{t}}) + (e ^{t}\ln q (q ^{e ^{t}})(p ^{e ^{t}})}{(1-q ^{e ^{t}})^{2}}\]

OpenStudy (m):

i'm guessing there is no short cut to solve this?

OpenStudy (m):

ln is not neccessary to derive this right

OpenStudy (dumbcow):

not really, the quotient rule is the shortcut i guess (f/g)' = f'g - fg'/g^2

OpenStudy (dumbcow):

i think it is, dervitive of a^x = ln(a)*a^x

OpenStudy (m):

ok thanks

OpenStudy (anonymous):

can i talk to anyone on fb?

OpenStudy (anonymous):

can i talk to anyone on fb?